OpenStudy (anonymous):

The function y= 3sin(2x) is: the inverse of a secant function undefined as x=2π even ( f(x) = f(-x) for all x) odd ( f(-x) = -f(-x) for all x) neither even nor odd

OpenStudy (anonymous):

Odd.

OpenStudy (roseflight):

\[\sin ^{-1}x \neq \sec x\] (sec x is the reciprocal of cosx not the inverse of sinx) I don't think sin x is undefined at 2pi... sinx is not even (\[\sin(-x) \neq \sin(x)\] for all x) However, sin IS odd (sin(-x) = -sin(-x) for all x)

OpenStudy (anonymous):

the function is odd?